For the reaction

NH4NO3(aq) ---->. N2O(g) + 2 H2O(l)

Delta G° = -183.7 kJ and

Delta H° = -149.6 kJ at 341 K and 1 atm.

This reaction is (reactant, product)_______favored under standard conditions at 341 K.

The entropy change for the reaction of 1.70 moles of NH4NO3(aq) at this temperature would be________ J/K.
Answer #1

1)

Since delta Go is negative, the reaction is product favoured.

Answer: product

2)

ΔGo = -183.7 KJ

ΔHo = -149.6 KJ

T = 341 K

use:

ΔGo = ΔHo - T*ΔSo

-183.7 = -149.6 - 341.0 *ΔSo

ΔSo = 0.100 KJ/K

ΔSo = 100 J/K

This is when 1 mol of NH4NO3 reacts as per given equation.

For 1.70 moles,

ΔSo = 1.70 * 100 J/K = 170 J/K

Answer: 170 J/K